OpenStudy (anonymous):

using cos(theta)=(-3/5) , pi<0<(3pi/2) , sin=.8, find cos(2theta)

OpenStudy (anonymous):

Sin = 0.8? or does sin = -0.8?

OpenStudy (anonymous):

I calculated .8

OpenStudy (anonymous):

sorry its sin(theta)=0.8

OpenStudy (anonymous):

Ah, watch what quadrant you are in. When you take a square root, it could be the positive answer, or the negative answer. In quadrant III, both sin and cos are negative.

OpenStudy (anonymous):

So, you know that cos(theta) = -3/5, and sin(theta) = -0.8. We want to know cos(2theta). What identity would we use?

OpenStudy (anonymous):

adding??

OpenStudy (anonymous):

lol i'm not sure.. my teacher isn't the best.

OpenStudy (anonymous):

\[\cos(2\Theta) = \cos^2(\Theta)-\sin^2(\Theta) = 2\cos^2(\Theta) - 1\]

OpenStudy (anonymous):

so, for 2cos^2(theta)-1, i'd plug in (-3/5)?

OpenStudy (anonymous):

Yeah. We'll check the sign of your result after you do that.

OpenStudy (anonymous):

ohkay one second

OpenStudy (anonymous):

how do i get cos^2 on my calculator

OpenStudy (anonymous):

You know cos(theta) = -3/5 So cos^2(theta) = (-3/5)^2 = 9/25 You can just put this value in for cos^2(theta)

OpenStudy (anonymous):

-.28 ?

OpenStudy (anonymous):

That's what I get :)

OpenStudy (anonymous):

ok thankyou so much. one more question?

OpenStudy (anonymous):

What quadrant do you expect to "land" in? Is the sign correct on your value for cos(2*theta)?

OpenStudy (anonymous):

I would expect to land in 2nd?

OpenStudy (anonymous):

because it says find sin(theta) which i did and found .8

OpenStudy (anonymous):

so , it'd be like (-2.8, .8)

OpenStudy (anonymous):

sin(theta) is -0.8, as we start in the third quadrant, where both cosine and sine are negative.

OpenStudy (anonymous):

why do we start in third quadrant?

OpenStudy (anonymous):

but yes, we expect to land in quadrant II, where cos is negative. We start in quadrant III because we are between pi and 3pi/2, which is quadrant III

OpenStudy (anonymous):

Oh because pi<0<-3pi/2 Ohkay ! yay ! haha thankyou.

OpenStudy (anonymous):

well can you help me find sin(2pi/3-theta)

OpenStudy (anonymous):

Since the absolute value of sin > absolute value of cos, we are closer to 3pi/2 than to pi. When we double the angle, we will land in quadrant II, cos is negative, and all is well.

OpenStudy (anonymous):

Well, the first thing we can do is to put that into an easier form. We know that sin(-x) = -sin(x). So, \[\sin(\frac{2 \pi}{3} - \Theta) = \sin(-(\Theta - \frac{2\pi}{3})) = -\sin(\Theta - \frac{2\pi}{3})\] Let me think of how to get a numerical value.

OpenStudy (anonymous):

So, you can do it with a calculator, if you are careful. But I can't come up with a good way to do it without one.

OpenStudy (anonymous):

its ohkay, i'm not really sure.. ohkay , so for -sin(theta-2pi/3) i can plug in -.8 for theta, and get the answer?

OpenStudy (anonymous):

No, -0.8 is the value for sin(theta). You'd need to find the actual angle. This is complicated slightly because calculators only report values for sin and cos in two quadrants. cos is reported in quadrants 1 and 2. sin in quadrants 1 and 4.

OpenStudy (anonymous):

ohhh im so confused ! haha ohkay thankyou.. could you help me with something else

OpenStudy (anonymous):

I can sure try :)

OpenStudy (anonymous):

this should be easier.. it says find all possible answers for sin(x)+cos(2x)=1

OpenStudy (anonymous):

Much easier :) We can use either the double angle formula or half angle formula to get all the trig functions to be of the same type. cos(2x) = 1-2sin^2(x) Therefore, sinx+1-2sin^2(x) = 1 Sin(x) - 2sin^2(x) = 0 Sin(x) = 2Sin^2(x) Sin(x) = 1/2

OpenStudy (anonymous):

ohkay, and i think this is multiplied sin*cos(2x)+cos*sin(2x)=0

OpenStudy (anonymous):

On the last one, make sure you find the values for x

OpenStudy (anonymous):

it doesn't give me values for x

OpenStudy (anonymous):

it just says to find possible answers..

OpenStudy (anonymous):

Right. The possible answers are all the values of x that make sin(x) equal to 1/2

OpenStudy (anonymous):

ohhhh , how do i get that

OpenStudy (anonymous):

Well, you can find one with a calculator easily. The other one is a related angle in another quadrant (sine is positive in quadrants I and II)

OpenStudy (anonymous):

For the last question you asked, we do something similar. Using two identities: \[Cos(2x) = 2Cos^2(x) - 1\] and \[Sin(2x) = 2Sin(x)Cos(x)\] We can rewrite the equation as: \[Sin(x)*(2Cos^2(x) - 1) + Cos(x)(2Cos(x)Sin(x)) = 0\] Which simplifies to: \[2Sin(x)Cos^2(x) - Sin(x) + 2Sin(x)Cos^2(x) = 0\] We add sin(x) to each side, and then divide by sin(x), as well as adding the other two terms together, and arrive at: \[4Cos^2(x) = 1\] \[Cos^2(x) = \frac{1}{4}\] \[Cos(x) = \frac{1}{2}\] Find the values for x that make this true.